OpenStudy (anonymous):

Driving toward home at a constant rate of 60 mph, you are 529 mi away. The distance, d, in miles away from home after t hours is given by d = 529 − 60t How far from home are you after 4 hours? Answer in units of mi

OpenStudy (anonymous):

d = 529 - 60t plug t = 4 d = 529 - 240 d = 289